# Floquet breathers in a modulated nonlinear lattice

### v3

@@ Abstract
We propose a space-time modulated system formed by an array of cantilevers where the on-site potential is provided by electromagnets fed with DC and AC currents. The modulation changes the phonon bands and the theory of breathers, or localized periodic solutions. We obtain breathers for many different frequencies, larger or smaller than the modulating one, but always commensurate. We analyze in particular the stability dependence on the frequency for equal-period, doubled.period and halved-period breathers. The results are of application for dynamic metamaterials.
@@ I Introduction
An experimental macroscopic array of cantilevers was introduced by Kimura and Hikihara in Ref. Kimura and Hikihara, 2009a.The system was driven by a periodic force applied to the support of the cantilevers, appearing as an independent external (additive) force in the dynamical equations. The restoring force had two contributions: one provided by the elastic properties of the cantilever, and other of magnetic nature, by fixing a magnet at the free end of the cantilever, and locating an electromagnet below it. The magnetic force due to permanent magnet-electromagnet interaction, can be tuned by changing an electric DC current, in this way, a tunable on-site potential was obtained.
@@
In this paper we present a variation of the model proposed in Ref. Kimura and Hikihara, 2009a, incorporating, besides the static (DC) current, an oscillating (AC) contribution in the electromagnets. This brings about an on-site potential that it is tunable and periodic in time. Furthermore, introducing phase differences between neighboring cantilevers, an on-site potential with a space-time modulation is achieved. An experimental and mathematical model with a linear on-site potential and nonlinear coupling has been studied,Chong et al. (2024) finding $q$-gap breathers. In our case the on-site potential is nonlinear and the coupling is linear.
The article is organized as follows: after the introduction in Sec. I, Sec. II presents the physical model and the derivation of the system of differential equations, both for the nonlinear system and its linear approximation corresponding to small oscillations. In Sec. III, the effect of space and time modulation on the dispersion bands is presented, while the theoretical deduction is located in Appendix C. Sec. IV presents changes brought about by modulation in breather theory, breather obtention, and linear stability. Sec. V, presents breathers with a period that is an integer of the modulation period, and thus with a frequency smaller than the modulating one, and their properties are analyzed. The case of the breather frequency as a multiple of the modulation frequency is presented in Sec. VI. Sec. VII considers also rational ratios of the breather and modulation frequencies and the possibility of the breathers to be in quadrature with the modulation. A systematic study of the evolution of linear stability with the frequency is performed. The final Sec. VIII presents the concluding remarks. Important but long explanations and mathematical derivations are included in three appendices in order not to interrupt the flow of the article focused on Floquet breathers. Appendix A deals with the concept of thermalization and numerical experiments that are used to check the theory. Appendix B present concepts and notation used in some of the main text and in Appendix C, where the mathematical derivations of the dispersion relations are included.
@@ II Physical system and dynamical equations
In this section we derive from physical laws the system dynamical equations. A sketch of the physical model is shown in Fig. 1.
Figure 1: (a) System of cantilevers mechanically coupled by a rod. (b) Side view of a cantilever. A permanent magnet is attached to the tip of the cantilever. An electromagnet is placed below the permanent magnet. (c) Definition of magnetic charges. Modified with permission from Ref. Kimura and Hikihara, 2009a
@@
Therefore the dynamical equations for an array of coupled cantilevers become:
@@
For small oscillations the on-site potential becomes also quadratic:
@@
The Hamiltonian equations $\dot{u}_{n}=-\frac{\displaystyle H}{\displaystyle\partial p_{n}}=p_{n}$ and $\dot{p}_{n}=-\frac{\displaystyle\partial H}{\displaystyle\partial u_{n}}$ are still valid and therefore the time derivative of the Hamiltonian becomes:
@@
In general this derivative will not be zero and also it will not be periodic in time, implying that $H$ will not be periodic either. For $h=0$, the coefficients of $\sin(\Omega t)$ have all the same negative sign, and $\textrm{d}H/\textrm{d}t=0$ only if all $u_{n}=0$. However $H$ can be conserved at periodic times. If $T_{\text{m}}=2\pi/\Omega$ is the modulation period and the coordinates $u_{n}$ are periodic with frequency $\omega_{\text{b}}$ and period $T_{\text{b}}=2\pi/\omega_{\text{b}}$, a sufficient condition for $H$ to be periodic with period $T_{\text{H}}$ is that there are integers $m_{\text{m}}$ and $m_{\text{b}}$ such that $T_{\text{H}}=m_{\text{m}}T_{\text{m}}=m_{\text{b}}T_{\text{b}}$ or, in terms of the frequencies:
@@ III Effect of space-time modulation on the phonon dispersion relation
In this section, we present the main consequences of space-time modulation of the linear system (7). They are also valid for small displacements $u_{n}(t)$ in the nonlinear model (9). Appendix A explains the concept of thermalizaton and numerical experiments to observe the phonon dispersion relation (PDR). The relatively detailed mathematical derivations based on Bloch theorem are included in the rest of App. C.
For $\delta=0$, that is, without modulation, the phonon dispersion relation is given by $\omega_{q}^{2}=\omega_{0}^{2}+2\kappa(1-\cos(q))$ as a function of the wave number $q$ of the phonon. In most of this work, we use the physical value of the coupling parameter $\kappa=0.0052$, with the phonon frequencies between $\omega_{0}=1$ and $\sqrt{\omega_{0}^{2}+4\kappa}=1.010$. The exception is this section, where for some simulations we choose a larger value $\kappa=0.1$, with the phonon band between the frequencies 1 and 1.183, in order to have a good visualization of the shape of the original phonon band and the new bands that appear with the modulation. The parameters are specified in the captions.
@@
Figure 3: Contour plot of the XTDFT after thermalization of the linear system with only time modulation. See App. C.1 for details. Parameters: $\kappa=0.1$, $\delta=0.075$, $h=0$, N=120 (Top panel) $\Omega=0.1$; (Bottom panel) $\Omega=0.01$.
@@
Figure 4: Contour plot of the XTDFT after thermalization of the linear system with both time and space modulation, together with the theoretical dispersion curves. See App. C.1 and C.4 for details. Parameters: $\delta=0.05$, $\Omega=0.05$, $\kappa=0.1$, $h=0.5$, N=120. Note that there is an area to the right where the harmonics are very close and become mixed.
@@ IV Breathers with space-time modulation
In this section, we study the conditions for breather existence in a modulated system, first reviewing the theory of exact breathers and adapting it to space-time modulation and to only time modulation.
@@ IV.A Exact traveling breathers
The resonant harmonics form straight lines within the $(q,\omega)$ space, called resonant lines. Their slope is $V_{b}$ and their intercept at the axis $q=0$ is $m\omega$, i.e, the moving frame frequency. Therefore, all the harmonics in a resonant line have the same moving frame frequency $m\omega_{F}$, with the integer $m$ indexing the lines.
@@ IV.B Application to space-time harmonically modulated systems
with $m_{\text{m}}$ an integer. This means that the modulating wave is a resonant harmonic with index $m_{\text{m}}$ and moving frame frequency $m_{\text{m}}\omega_{F}$ and therefore the breather and modulating waves have commensurate moving frame frequencies.
Lemma 1 A necessary condition for an exact breather within a space-time modulated system with harmonic modulation is that the modulating harmonic is within a resonant line with the breather, or, in other words, that their moving frame frequencies are both integer multiples of the fundamental frequency $\omega_{F}$. Therefore $\omega_{b}$ and $\Omega$ are commensurate.
Note, that an exact moving breather is not periodic in time and has not a definite frequency, this condition applies to the moving frame frequencies and it is therefore different from Eq. 15, which reappears below.
@@ IV.C Breather obtention in the time-modulated system
The most direct way is working in the real space with the Newton and shooting method. Testing some different localized initial conditions and observing an approximate periodical behavior with period close to $T_{\text{b}}$, the intended one. The initial conditions are used as a seed for the Newton method, in order to obtain the initial conditions corresponding to the exact $T_{\text{b}}$-periodic behavior. Exact means with a precision of $10^{-15}$, that is, very close to the computer machine precision. The Newton method, a numerical application of the implicit function theorem, can be used because the system has no time invariance due to the time-periodic term in the potential, and a solution is unique in its neighborhood. The precision can be refined further by working in the space of the frequencies. If we consider time-reversible solutions, they can be written as the inverse discrete Fourier transform $u_{n}(t)=\sum_{k=-k_{\text{m}}}^{k_{\text{m}}}z_{k,n}\exp(\textrm{i}k\omega_{\text{b}}t)$. Due to $u_{n}(t)$ being real and time-symmetric, $z_{k,n}^{*}=z_{-k,n}$ and $z_{k,n}=z_{-k,n}$ and real. Then, $u_{n}(t)=z_{0,n}+\sum_{k=1}^{k_{\text{m}}}2z_{k,n}\cos(k\omega_{\text{b}}t)$, and the Newton method can be coded to obtain the $k_{\text{m}}+1$ real coefficients $z_{k,n}$ for each site $n$. The value of $k_{\text{m}}=15$ provides excellent results.
@@ IV.D Stability of breathers in the time-modulated system
It can also be checked numerically that the phase mode $\dot{u}_{\text{b}}(0),\ddot{u}_{\text{b}}(0)]^{t}=[0,\ddot{u}_{\text{b}}(0)]^{t}$ if not an eigenvector of the Floquet matrix $F$ with other eigenvalue. The isolated eigenvalues outside the phonon band correspond to eigenvectors with the same localization of the breather that are internal modes of the breather. In particular, when they have the same symmetry of the breather, they are called breathing modes.Johansson and Aubry (2000) Simulations with the eigenvectors added as a perturbation to the breather show that they produce a beating of the amplitude of the breather which justify their name.
@@ IV.E Extended autonomous symplectic system
The Hamiltonian structure of the dynamical equations, i.e., $\dot{u}_{n}=\partial H/\partial p_{n}$, $\dot{p}_{n}=-\partial H/\partial u_{n}$ is kept for the conjugate pairs of variables $u_{n}$, $p_{n}$, although the system is no longer autonomous. By defining $s=t$ or to be precise as (23) is $T_{\text{b}}$-periodic, as $s=\mathrm{rem}(t,T_{\text{b}})$, the remainder of $t/T_{\text{b}}$, i.e., $\mathrm{rem}(nT_{\text{b}}+t,T_{\text{b}})=t$ and $\mathrm{rem}(-nT_{\text{b}}-t,T_{\text{b}})=-t$, both for $n>0$ and $t>0$, so as $s$ is $T_{\text{b}}$ periodic. Then $\dot{s}=1$ as usual, and the system with the extra variable has become autonomous, but not symplectic as $s$ has no conjugate variable yet. The extra variable $h$ to make the system symplectic is obtained by defining a new Hamiltonian $K=H+h$, with $h$ the conjugate momentum of $s$ as $\dot{s}=\partial K/\partial h=1$ as previously obtained. Then, $\dot{h}=-\partial K/\partial s=-\partial H/\partial s=-\partial H/\partial t=-\textrm{d}H/\textrm{d}t$, or $h=E_{0}-H$, with $E_{0}$ the initial energy at $t=0$.Marthinsen and Owren (2016) The extended system is now symplectic and $\dot{u}_{b},\dot{s}]$ is now solution of (24) and being $T_{\text{b}}$ periodic, the Floquet multiplier +1+1 corresponding to the phase mode reappears. There is an extra +1+1 multiplier, because, symplectiness, implies that the multipliers are in pairs $\lambda,1/\lambda$, the extra mode corresponding to a solution with slightly different energy or growth mode.
@@ V Breathers with time modulation and $\Omega=n_{T}\omega_{\text{b}}$
We consider the nonlinear dynamical system (9) initially for the cases $\omega_{\text{b}}=\Omega/n_{T}$ ($T_{\text{b}}=n_{T}T_{\text{m}}$), and $\omega_{\text{b}}=n_{T}\Omega$ ($T_{\text{b}}=T_{\text{m}}/n_{T}$), $n_{T}$ an integer. The first one corresponds to the well known phenomenon of period doubling, tripling, etc., while the latter correspond to a breather that oscillates several times faster than the modulation. According to the deduction in Sec. IV, fractional frequencies are also possible and we will consider them in Sec. VII for different breather frequencies.
There is no dissipation of energy, but as the system is parametrically forced the energy is conserved only at discrete times separated by the larger of the two periods, the breather period $T_{\text{b}}$ and the modulating one $T_{m}$. This is shown in Fig. 5 for $T_{\text{b}}=2T_{m}$.
In this section we consider $\omega_{\text{b}}=\Omega/n_{T}$ and leave the condition of $\omega_{\text{b}}>\Omega$ for the next section.
The dynamical equations (9) are invariant under the transformation $t\rightarrow t+T_{m}$, therefore, if a phonon $(q,\omega)$, i.e., $u=\exp(\textrm{i}[qn-\omega t])$ is a solution of the linearized equation, then also $\exp(\textrm{i}m\Omega t)u$ is also a solution, or, in other words, there is another phonon band with frequencies $\omega^{\prime}=\omega+m\Omega$, with $m$ a positive or negative integer. But there are also two phonon bands, with positive and negative frequencies. Then, if there is a phonon $(q,\omega)$, also there exists the phonon $(q,\Omega-\omega)$. As the unmodulated phonon band for $\kappa=0.052$ is in $\omega_{0}=1,\omega_{\text{max}}=1.01]$, for $\Omega\simeq 2$, a new phonon band appears in $\Omega-\omega_{\text{max}}\simeq 0.98,\Omega-\omega_{0}\simeq 1]$. The phonon bands in the modulated system change position, and $\Omega=2\omega_{\text{b}}=1.96$, but those bands are very close with the possibility of interfering, depending on the exact value of $\omega_{\text{b}}$ and actual modulated phonons. This problem does not exist for $\Omega\geq 3$ as the band $\Omega-\omega$, for $\omega$ in the unmodulated phonon band is much more separated.
For $\Omega=2\omega_{\text{b}}$, and changing slightly $\omega_{\text{b}}$, there is an interval where the two phonon bands coincide and then the system becomes stimulated by $\Omega$ and the variables diverge, as it is also shown in Sec. VII. A dissipation term can be added, but as this is a very special case, we prefer not to modify the system and simply avoid those combination of frequencies.

### v4

@@ Abstract
We propose a space-time modulated system formed by an array of cantilevers where the on-site potential is provided by electromagnets fed with DC and AC currents. The modulation changes the phonon bands and the theory of breathers, or localized periodic solutions. We obtain breathers for many different frequencies, larger or smaller than the modulating one, but always commensurate. We analyze in particular the stability dependence on the frequency for equal-period, doubled-period and halved-period breathers. The results are of application for dynamic metamaterials.
@@ I Introduction
An experimental macroscopic array of cantilevers was introduced by Kimura and Hikihara in Ref. Kimura and Hikihara, 2009a. The system was driven by a periodic force applied to the support of the cantilevers, appearing as an independent external (additive) force in the dynamical equations. The restoring force had two contributions: one provided by the elastic properties of the cantilever, and the other of magnetic nature, by fixing a magnet at the free end of the cantilever, and locating an electromagnet below it. The magnetic force due to permanent magnet-electromagnet interaction, can be tuned by changing an electric DC current, in this way, a tunable on-site potential was obtained.
@@
In this paper we present a variation of the model proposed in Ref. Kimura and Hikihara, 2009a, incorporating, besides the static (DC) current, an oscillating (AC) contribution in the electromagnets. This brings about an on-site potential that it is tunable and periodic in time. Furthermore, introducing phase differences between neighboring cantilevers, an on-site potential with a space-time modulation is achieved. An experimental and mathematical model with a linear on-site potential and nonlinear coupling has been studied,Chong et al. (2024) finding $q$-gap breathers. In our case, the on-site potential is nonlinear and the coupling is linear.
The article is organized as follows: after the introduction in Sec. I, Sec. II presents the physical model and the derivation of the system of differential equations, both for the nonlinear system and its linear approximation corresponding to small oscillations. In Sec. III, the effect of space and time modulation on the dispersion bands is presented, while the theoretical deduction is located in Appendix C. Sec. IV presents changes brought about by modulation in breather theory, breather obtention, and linear stability. Sec. V, presents breathers with a period that is an integer of the modulation period, and thus with a frequency smaller than the modulating one, and their properties are analyzed. The case of the breather frequency as a multiple of the modulation frequency is presented in Sec. VI. Sec. VII considers also rational ratios of the breather and modulation frequencies, and the possibility of the breathers to be in quadrature with the modulation. A systematic study of the evolution of linear stability with the frequency is performed. The final Sec. VIII presents the concluding remarks. Important but long explanations and mathematical derivations are included in three appendices in order not to interrupt the flow of the article focused on Floquet breathers. Appendix A deals with the concept of thermalization and numerical experiments that are used to check the theory. Appendix B presents concepts and notation used in some of the main text and in Appendix C, where the mathematical derivations of the dispersion relations are included.
@@ II Physical system and dynamical equations
In this section, we derive from physical laws the system dynamical equations. A sketch of the physical model is shown in Fig. 1.
Figure 1: (a) System of cantilevers mechanically coupled by a rod. (b) Side view of a cantilever. A permanent magnet is attached to the tip of the cantilever. An electromagnet is placed below the permanent magnet. (c) Definition of magnetic charges.
@@
Therefore, the dynamical equations for an array of coupled cantilevers become:
@@
For small oscillations, the on-site potential becomes also quadratic:
@@
The Hamiltonian equations $\dot{u}_{n}=-\frac{\displaystyle H}{\displaystyle\partial p_{n}}=p_{n}$ and $\dot{p}_{n}=-\frac{\displaystyle\partial H}{\displaystyle\partial u_{n}}$ are still valid and, therefore, the time derivative of the Hamiltonian becomes:
@@
In general, this derivative will not be zero and also it will not be periodic in time, implying that $H$ will not be periodic either. For $h=0$, the coefficients of $\sin(\Omega t)$ have all the same negative sign, and $\textrm{d}H/\textrm{d}t=0$ only if all $u_{n}=0$. However, $H$ can be conserved at periodic times. If $T_{\text{m}}=2\pi/\Omega$ is the modulation period and the coordinates $u_{n}$ are periodic with frequency $\omega_{\text{b}}$ and period $T_{\text{b}}=2\pi/\omega_{\text{b}}$, a sufficient condition for $H$ to be periodic with period $T_{\text{H}}$ is that there are integers $m_{\text{m}}$ and $m_{\text{b}}$ such that $T_{\text{H}}=m_{\text{m}}T_{\text{m}}=m_{\text{b}}T_{\text{b}}$ or, in terms of the frequencies:
@@ III Effect of space-time modulation on the phonon dispersion relation
In this section, we present the main consequences of space-time modulation of the linear system (7). They are also valid for small displacements $u_{n}(t)$ in the nonlinear model (9). Appendix A explains the concept of thermalization and numerical experiments to observe the phonon dispersion relation (PDR). The relatively detailed mathematical derivations based on Bloch theorem are included in the rest of App. C.
For $\delta=0$, that is, without modulation, the phonon dispersion relation is given by $\omega_{q}^{2}=\omega_{0}^{2}+2\kappa(1-\cos(q))$ as a function of the wave number $q$ of the phonon. In most of this work, we use the physical value of the coupling parameter $\kappa=0.0052$, with the phonon frequencies between $\omega_{0}=1$ and $\sqrt{\omega_{0}^{2}+4\kappa}=1.010$. The exception is this section, where for some simulations we choose a larger value $\kappa=0.10$, with the phonon band between the frequencies 1 and 1.183, in order to have a good visualization of the shape of the original phonon band and the new bands that appear with the modulation. The parameters are specified in the captions.
@@
Figure 3: Contour plot of the XTDFT after thermalization of the linear system with only time modulation. See App. C.1 for details. Parameters: $\kappa=0.10$, $\delta=0.075$, $h=0$, N=120 (Top panel) $\Omega=0.1$; (Bottom panel) $\Omega=0.01$.
@@
Figure 4: Contour plot of the XTDFT after thermalization of the linear system with both time and space modulation, together with the theoretical dispersion curves. See App. C.1 and C.4 for details. Parameters: $\delta=0.05$, $\Omega=0.05$, $\kappa=0.10$, $h=0.5$, N=120. Note that there is an area to the right where the harmonics are very close and become mixed.
@@ IV Breathers with space-time modulation
In this section, we study the conditions for breather existence in a modulated system, first reviewing the theory of exact moving breathers and adapting it to space-time modulation and to only time modulation.
@@ IV.A Exact traveling breathers
The resonant harmonics form straight lines within the $(q,\omega)$ space, called resonant lines. Their slope is $V_{b}$ and their intercept at the axis $q=0$ is $m\omega$, i.e., the moving frame frequency. Therefore, all the harmonics in a resonant line have the same moving frame frequency $m\omega_{F}$, with the integer $m$ indexing the lines.
@@ IV.B Application to space-time harmonically modulated systems
with $m_{\text{m}}$ an integer. This means that the modulating wave is a resonant harmonic with index $m_{\text{m}}$ and moving frame frequency $m_{\text{m}}\omega_{F}$ and, therefore, the breather and modulating waves have commensurate moving frame frequencies.
Lemma 1 A necessary condition for an exact breather within a space-time modulated system with harmonic modulation is that the modulating harmonic is within a resonant line with the breather, or, in other words, that their moving frame frequencies are both integer multiples of the fundamental frequency $\omega_{F}$. Therefore, $\omega_{b}$ and $\Omega$ are commensurate.
Note, that an exact moving breather is not periodic in time and has not a definite frequency, this condition applies to the moving frame frequencies, and it is, therefore, different from Eq. 15, which reappears below.
@@ IV.C Breather obtention in the time-modulated system
The most direct way is working in the real space with the Newton and shooting method. Testing some different localized initial conditions and observing an approximate periodical behavior with a period close to $T_{\text{b}}$, the intended one. The initial conditions are used as a seed for the Newton method, in order to obtain the initial conditions corresponding to the exact $T_{\text{b}}$-periodic behavior. Exact means with a precision of $10^{-15}$, that is, very close to the computer machine precision. The Newton method, a numerical application of the implicit function theorem, can be used because the system has no time invariance due to the time-periodic term in the potential, and a solution is unique in its neighborhood. The precision can be refined further by working in the space of the frequencies. If we consider time-reversible solutions, they can be written as the inverse discrete Fourier transform $u_{n}(t)=\sum_{k=-k_{\text{m}}}^{k_{\text{m}}}z_{k,n}\exp(\textrm{i}k\omega_{\text{b}}t)$. Due to $u_{n}(t)$ being real and time-symmetric, $z_{k,n}^{*}=z_{-k,n}$ and $z_{k,n}=z_{-k,n}$ and real. Then, $u_{n}(t)=z_{0,n}+\sum_{k=1}^{k_{\text{m}}}2z_{k,n}\cos(k\omega_{\text{b}}t)$, and the Newton method can be coded to obtain the $k_{\text{m}}+1$ real coefficients $z_{k,n}$ for each site $n$. The value of $k_{\text{m}}=15$ provides excellent results.
@@ IV.D Stability of breathers in the time-modulated system
It can also be checked numerically that the phase mode $\dot{u}_{\text{b}}(0),\ddot{u}_{\text{b}}(0)]^{t}=[0,\ddot{u}_{\text{b}}(0)]^{t}$ is not an eigenvector of the Floquet matrix $F$ with another eigenvalue. The isolated eigenvalues outside the phonon band correspond to eigenvectors with the same localization of the breather, that are internal modes of the breather. In particular, when they have the same symmetry of the breather, they are called breathing modes.Johansson and Aubry (2000) Simulations with the eigenvectors added as a perturbation to the breather show that they produce a beating of the amplitude of the breather which justify their name.
@@ IV.E Extended autonomous symplectic system
The Hamiltonian structure of the dynamical equations, i.e., $\dot{u}_{n}=\partial H/\partial p_{n}$, $\dot{p}_{n}=-\partial H/\partial u_{n}$ is kept for the conjugate pairs of variables $u_{n}$, $p_{n}$, although the system is no longer autonomous. By defining $s=t$ or to be precise as (23) is $T_{\text{b}}$-periodic, as $s=\mathrm{rem}(t,T_{\text{b}})$, the remainder of $t/T_{\text{b}}$, i.e., $\mathrm{rem}(nT_{\text{b}}+t,T_{\text{b}})=t$ and $\mathrm{rem}(-nT_{\text{b}}-t,T_{\text{b}})=-t$, both for $n>0$ and $t>0$, so as $s$ is $T_{\text{b}}$ periodic. Then $\dot{s}=1$ as usual, and the system with the extra variable has become autonomous, but not symplectic as $s$ has no conjugate variable yet. The extra variable $h$ to make the system symplectic is obtained by defining a new Hamiltonian $K=H+h$, with $h$ the conjugate momentum of $s$ as $\dot{s}=\partial K/\partial h=1$ as previously obtained. Then, $\dot{h}=-\partial K/\partial s=-\partial H/\partial s=-\partial H/\partial t=-\textrm{d}H/\textrm{d}t$, or $h=E_{0}-H$, with $E_{0}$ the initial energy at $t=0$.Marthinsen and Owren (2016) The extended system is now symplectic and $\dot{u}_{b},\dot{s}]$ is now solution of (24) and being $T_{\text{b}}$ periodic, the Floquet multiplier +1+1 corresponding to the phase mode reappears. There is an extra +1+1 multiplier, because symplectiness implies that the multipliers are in pairs $\lambda,1/\lambda$, the extra mode corresponding to a solution with slightly different energy or growth mode.
@@ V Breathers with time modulation and $\Omega=n_{T}\omega_{\text{b}}$
We consider the nonlinear dynamical system (9) initially for the cases $\omega_{\text{b}}=\Omega/n_{T}$ ($T_{\text{b}}=n_{T}T_{\text{m}}$), and $\omega_{\text{b}}=n_{T}\Omega$ ($T_{\text{b}}=T_{\text{m}}/n_{T}$), $n_{T}$ an integer. The first one corresponds to the well known phenomenon of period doubling, tripling, etc., while the latter correspond to a breather that oscillates several times faster than the modulation. According to the deduction in Sec. IV, fractional frequencies are also possible, and we will consider them in Sec. VII for different breather frequencies.
There is no dissipation of energy, but as the system is parametrically forced, the energy is conserved only at discrete times separated by the larger of the two periods, the breather period $T_{\text{b}}$ and the modulating one $T_{m}$. This is shown in Fig. 5 for $T_{\text{b}}=2T_{m}$.
In this section, we consider $\omega_{\text{b}}=\Omega/n_{T}$ and leave the condition of $\omega_{\text{b}}>\Omega$ for the next section.
The dynamical equations (9) are invariant under the transformation $t\rightarrow t+T_{m}$, therefore, if a phonon $(q,\omega)$, i.e., $u=\exp(\textrm{i}[qn-\omega t])$ is a solution of the linearized equation, then also $\exp(\textrm{i}m\Omega t)u$ is also a solution, or, in other words, there is another phonon band with frequencies $\omega^{\prime}=\omega+m\Omega$, with $m$ a positive or negative integer. But there are also two phonon bands, with positive and negative frequencies. Then, if there is a phonon $(q,\omega)$, there also exists the phonon $(q,\Omega-\omega)$. As the unmodulated phonon band for $\kappa=0.052$ is in $\omega_{0}=1,\omega_{\text{max}}=1.01]$, for $\Omega\simeq 2$, a new phonon band appears in $\Omega-\omega_{\text{max}}\simeq 0.98,\Omega-\omega_{0}\simeq 1]$. The phonon bands in the modulated system change position, and $\Omega=2\omega_{\text{b}}=1.96$, but those bands are very close with the possibility of interfering, depending on the exact value of $\omega_{\text{b}}$ and actual modulated phonons. This problem does not exist for $\Omega\geq 3$ as the band $\Omega-\omega$, for $\omega$ in the unmodulated phonon band is much more separated.
For $\Omega=2\omega_{\text{b}}$, and changing slightly $\omega_{\text{b}}$, there is an interval where the two phonon bands coincide and then the system becomes stimulated by $\Omega$ and the variables diverge, as it is also shown in Sec. VII. A dissipation term can be added, but as this is a very special case, we prefer not to modify the system and simply avoid those combinationsof frequencies.